# SplatSim: Zero-Shot Sim2Real Transfer of RGB Manipulation Policies Using Gaussian Splatting

### v1

@@ Abstract
Sim2Real transfer, particularly for manipulation policies relying on RGB images, remains a critical challenge in robotics due to the significant domain shift between synthetic and real-world visual data. In this paper, we propose SplatSim, a novel framework that leverages Gaussian Splatting as the primary rendering primitive to reduce the Sim2Real gap for RGB-based manipulation policies. By replacing traditional mesh representations with Gaussian Splats in simulators, SplatSim produces highly photorealistic synthetic data while maintaining the scalability and cost-efficiency of simulation. We demonstrate the effectiveness of our framework by training manipulation policies within SplatSim and deploying them in the real world in a zero-shot manner, achieving an average success rate of 86.25%, compared to 97.5% for policies trained on real-world data.
@@ I INTRODUCTION
The Sim2Real problem, a focal challenge in robotics, pertains to the transfer of control policies learned in simulated environments to real world settings. Recently, significant progress has been made in deploying controllers trained in simulation to the real world in a zero-shot manner. Robots have demonstrated the ability to walk on rough terrains [1, 2], perform in-hand object rotation [3, 4, 5, 6], and grasp previously unseen objects [7]. Notably, all of these methods rely on perception modalities like depth, tactile sensing, or point cloud inputs, which have gained significant attention due to the relatively small Sim2Real gap they offer. The reduced discrepancy between simulated and real-world data in these modalities has led to remarkable progress, reinforcing the idea that modalities that can be simulated well, can be transferred well.
@@
  • We show how to leverage Robot Splat Models and Object Splat Models, along with the simulator as a physics backend, to generate photorealistic trajectories of robot-object interactions. Our method eliminates the need for real world data collection to learn these interactions, and relies solely on an initial video of the scene with the robot. We further demonstrate how these renderings, when combined with simulated demonstrations, can be utilized to generate high-quality synthetic datasets for behavior cloning methods.
  • We demonstrate the effectiveness of our framework by deploying RGB policies, trained entirely in simulation, to the real world in a zero-shot manner across four tasks, achieving an average success rate of 86.25%, compared to 97.5% for policies trained on real-world data.
@@ Acknowledgement
This work is in part supported by NSF/USDA-NIFA AIIRA AI Research Institute 2021-67021-35329 and USDA-NIFA/NSF National Robotics Initiative 2021-67021-35974.

### v2

@@ Abstract
Sim2Real transfer, particularly for manipulation policies relying on RGB images, remains a critical challenge in robotics due to the significant domain shift between synthetic and real-world visual data. In this paper, we propose SplatSim, a novel framework that leverages Gaussian Splatting as the primary rendering primitive to reduce the Sim2Real gap for RGB-based manipulation policies. By replacing traditional mesh representations with Gaussian Splats in simulators, SplatSim produces highly photorealistic synthetic data while maintaining the scalability and cost-efficiency of simulation. We demonstrate the effectiveness of our framework by training manipulation policies within SplatSim and deploying them in the real world in a zero-shot manner, achieving an average success rate of 86.25%, compared to 97.5% for policies trained on real-world data. Videos can be found on our project page: https://splatsim.github.io
@@ I INTRODUCTION
The Sim2Real problem, a focal challenge in robotics, pertains to the transfer of control policies learned in simulated environments to real world settings. Recently, significant progress has been made in deploying controllers trained in simulation to the real world in a zero-shot manner. Robots have demonstrated the ability to walk on rough terrains [1, 2], perform in-hand object rotation [3, 4, 5, 6], and grasp previously unseen objects [7]. Notably, all of these methods rely on perception modalities like depth, tactile sensing, or point cloud inputs, which have gained significant attention due to the relatively small Sim2Real gap they offer. The reduced discrepancy between simulated and real-world data in these modalities has led to remarkable progress, reinforcing the idea that policies trained on modalities that can be simulated well, can be transferred well.
@@
  • We show how to leverage Robot Splat Models and Object Splat Models, along with the simulator as a physics backend, to generate photorealistic trajectories of robot-object interactions. Our method eliminates the need for the real-world data collection to learn these interactions, and relies solely on an initial video of the static scene with the robot. We further demonstrate how these renderings, when combined with simulated demonstrations, can be utilized to generate high-quality synthetic datasets for behavior cloning methods.
  • We demonstrate the effectiveness of our framework by deploying RGB policies, trained entirely in simulation, to the real world in a zero-shot manner across four tasks, achieving an average success rate of 86.25%, compared to 97.5% for policies trained on the real-world data.
@@ Acknowledgement
We would like to express our gratitude to Prof. Shubham Tulsiani for his valuable insights during the early development of this idea. We also extend our thanks to Anurag Ghosh and Ayush Jain for their thoughtful feedback in refining the manuscript. This work is in part supported by NSF/USDA-NIFA AIIRA AI Research Institute 2021-67021-35329 and USDA-NIFA/NSF National Robotics Initiative 2021-67021-35974.